# Statistical modeling of SOFR term structure

### v2

@@ 4.4 The forward curve model
We model the daily values of the 9-dimensional process $x=(x_{t})_{t=0}^{\infty}$ defined in (15) and (16) by another vector autoregressive model of the form (17). The vector autoregressive model is fitted to daily data constructed from the historical values of the target rate $L$ and the daily closing prices of futures contracts as described in Section 4.2. The parameter matrix and the distribution of the innovations are estimated as described in Section 4.3 in the context of the macroeconomic variables. However, to simplify the statistical analysis and the modelling, we restrict the autoregressive matrix to be diagonal. The estimated coefficients are given in Table 9. All diagonal values are negative which indicates mean reversion and stationarity in the components of $x$. Table 10 reports the correlations and the variances of the residual term.
@@ 4.5 Incorporation of user’s views
We apply the calibration procedure from Section 4.2 of [aap21], which allows us to incorporate both short term forecasts as well as long-term views in both the macroeconomic model as well as the forward curve model described in Sections 4.3 and 4.4 above. The calibration is achieved by choosing the sequence of vectors $a_{t}$ in (17) so that the values of $x_{t}$ have desired median values at given points in time. Since both the macro model and the forward curve model are stationary, the model calibration according to Section 4.2 of [aap21] is particularly easy.
@@ 5.2 Simulating SOFR derivatives payouts
If exercised, a swaption, with strike $X$ and maturity $T$, converts to a short position in an OIS with swap rate $X$ and $K$ yearly payments; see [CMEIRprod, CMEswaption]. Assuming perfectly liquid zero-coupon bonds, the market price of such an OIS would be
@@ 5.3 Indifference pricing of SOFR derivatives
A common approach to derivative pricing is to construct a market model together with a martingale measure which is then used for pricing derivatives by taking expectations of their cashflows. Fixing a pricing measure, however, leads to a linear pricing rule, which contradicts the nonlinearities observed in practice. Such problems are addressed explicitly by the techniques of incomplete markets, where one first develops a term structure model under the “statistical” measure and then defines prices according to the indifference principle; see e.g. [car9], [pen14] and their references. This section uses the model developed in Sections 2,  3, and 4 in indifference pricing of the derivatives described in the previous section. The exercise below is quite simple in that the only “hedging instrument” is the money market account represented by the overnight Repo market. More sophisticated hedging strategies will be built in the follow-up paper that takes the stochastic model developed in this paper as the description of the underlying risk factors.
Let $c$ be the payout at maturity $t_{1}$ of a SOFR derivative and $U:\mathbbm{R}\to\mathbbm{R}$ a concave strictly increasing utility function. The utility indifference selling price at $t_{0}$ is the least amount of cash $\alpha$ for which an agent, with initial wealth $\bar{w}$, is indifferent between selling the claim $c$ for $\alpha$, and not doing the trade. In the case of the futures and swaption, where the price is paid at maturity, the indifference price is the solution $\alpha$ to the equation
@@
Assuming that the futures options are only exercised at maturity, their indifference prices are given by the same formula.
The indifference price $\alpha$ in (19) can be computed by numerically evaluating the expectations. In the numerical computations below, we will use Monte Carlo with antithetic sampling. That is, we simulate a random sample of $N$ scenarios and obtain another $N$ scenarios by reflecting the underlying Gaussian process.
Figures 12 and 13 plot the indifference selling prices for the four SOFR derivatives studied in Section 5.2. Each plot in Figure 12 displays the indifference price of the swaption and the futures as a function of the initial wealth $\bar{w}$ and the risk aversion parameter $\rho$. As $\bar{w}$ increases, the prices increase. With an initial financial position of cash only, the agent is less desperate for extra cash and asks for a higher price to compensate for taking on the additional risk of selling the derivative. As $\rho$ increases, the prices increase. Because their initial financial position consists only of cash, the higher the value of $\rho$ the more risk-averse the agent is. They require more money to sell the derivative. This is also observed in the plots of Figure 13 which display the options prices as functions of the risk aversion parameter $\rho$.
Figure 12: Selling price of SOFR three-month futures and swaption as a function of the initial wealth $\bar{w}$ and the risk aversion parameter $\rho$ (scaled by a factor of $10^{-3}$) as of 28 August 2024.
Figure 13: Selling price of SOFR options on three-month futures as a function of the risk aversion parameter $\rho$ (scaled by a factor of $10^{-3}$) as of 28 August 2024.
@@ 6 Arbitrage
The following theorem is from [js98, Theorem 3], but one should note that part g of the statement there has a typo: the symbol $Q_{n}$ should be replaced by $\bar{Q}_{n}$. An alternative proof and an extended statement can be found in [pp24, Remark 2.120]. The ${\cal F}_{t}$}-conditional support of a random variable $X\in L^{0}(\Omega,{\cal F},P;\mathbbm{R}^{d})$ is the smallest closed-valued ${\cal F}_{t}$-measurable mapping $\mathop{\rm supp}\nolimits_{t}X:\Omega\rightrightarrows\mathbbm{R}^{d}$ such that $X\in\mathop{\rm supp}\nolimits_{t}X$ almost surely. The convex hull and the relative interior of a set $C\subseteq\mathbbm{R}^{d}$ will be denoted by $\mathop{\rm co}C$ and $\mathop{\rm ri}C$, respectively.
@@ 6.1 Arbitrage in zero-coupon bonds
By [pp24, Theorem 2.43.4], the continuity of $H$ implies
@@
The claim now follows from the fact that $\mathop{\rm ri}\mathop{\rm co}\mathop{\rm cl}C=\mathop{\rm ri}\mathop{\rm co}C$ for any set $C$ in a Euclidean space; see [roc70a, Theorem 6.3]. ∎
@@
i.e. if the conditional support of the day-ahead forward curve contains every positive forward curve. This condition is far from necessary but it could be enforced e.g. by multiplying each $F_{t}(s)$ by $e^{\epsilon_{t}}$, where $\epsilon_{t}$ are independent Gaussian variables with nonzero variance. Such additional randomness could be interpreted as observation errors, as is common in the theory of affine term structure models; see e.g. [cheridito2007market, Section 4]. Taking the variance of $\epsilon_{t}$ sufficiently close to zero, the additional randomness would not have a material effect on portfolio optimization or indifference pricing in the face of bid-ask spreads. More importantly, one cannot trade perfectly liquid zero-coupon bonds with arbitrary maturities in practice, so the no-arbitrage conditions derived above aren’t all that relevant in real-life applications.
@@ 6.2 Arbitrage in futures market
By [pp24, Theorem 2.43.4], continuity of the exponential function implies
@@
The claim now follows from the fact that $\mathop{\rm ri}\mathop{\rm co}\mathop{\rm cl}C=\mathop{\rm ri}\mathop{\rm co}C$ for any set $C$ in a Euclidean space; see [roc70a, Theorem 6.3]. ∎

### v3

@@ 4.4 The forward curve model
We model the daily values of the 9-dimensional process $x=(x_{t})_{t=0}^{\infty}$ defined in (17) and (18) by another vector autoregressive model of the form
$\Delta x_{t}=Ax_{t-1}+a_{t}+\epsilon_{t},$ (21)
where $A\in\mathbbm{R}^{9\times 9}$ and $a_{t}\in\mathbbm{R}^{9}$ are given parameters and $\epsilon_{t}$ are iid Gaussian vectors in $\mathbbm{R}^{9}$ with zero mean and covariance matrix $\Sigma$. The vector autoregressive model is fitted to daily data constructed from the historical values of the target rate $L$ and the daily closing prices of futures contracts as described in Section 4.2. The parameter matrix and the distribution of the innovations are estimated as described in Section 4.3 in the context of the macroeconomic variables. However, to simplify the statistical analysis and the modelling, we restrict the autoregressive matrix to be diagonal. One could, of course, allow for general AR-matrices but the interpretation of the off-diagonal coefficients would not be quite straightforward. The diagonal elements, on the other hand, correspond to mean reversion which is a natural feature for the shape of the modelled forward curve. The estimated coefficients are given in Table 9. All diagonal values are negative which indicates mean reversion and stationarity in the components of $x$. Table 10 reports the correlations and the variances of the residual term.
@@ 4.5 Incorporation of user’s views
We apply the calibration procedure from Section 4.2 of [27], which allows us to incorporate both short-term forecasts as well as long-term views in both the macroeconomic model as well as the forward curve model described in Sections 4.3 and 4.4 above. The calibration is achieved by choosing the sequence of vectors $a_{t}$ in (20) so that the values of $x_{t}$ have desired median values at given points in time. Since both the macro model and the forward curve model are stationary, the model calibration according to Section 4.2 of [27] is particularly easy.
@@ 5.2 Simulating SOFR derivatives payouts
If exercised, a swaption, with strike $X$ and maturity $T$, converts to a short position in an OIS with swap rate $X$ and $K$ yearly payments; see [15, 16]. Assuming perfectly liquid zero-coupon bonds, the market price of such an OIS would be
@@ 5.3 Indifference pricing of SOFR derivatives
A common approach to derivative pricing is to construct a market model together with a martingale measure which is then used for pricing derivatives by taking expectations of their cashflows. Fixing a pricing measure, however, leads to a linear pricing rule, which contradicts the nonlinearities observed in practice. Such problems are addressed explicitly by the techniques of incomplete markets, where one first develops a term structure model under the “statistical” measure and then defines prices according to the indifference principle; see e.g. [9], [34] and their references. This section uses the model developed in Sections 2,  3, and 4 in indifference pricing of the derivatives described in the previous section. The exercise below is quite simple in that the only “hedging instrument” is the money market account represented by the overnight Repo market. More sophisticated hedging strategies will be built in a follow-up paper that uses exchange-traded derivatives to hedge OTC options.
Let $c$ be the payout at maturity $t_{1}$ of a SOFR derivative and $U:\mathbbm{R}\to\mathbbm{R}$ a concave strictly increasing utility function. The utility indifference selling price at $t_{0}$ is the least amount of cash $\alpha$ for which an agent, with initial wealth $\bar{w}$, is indifferent between selling the claim $c$ for $\alpha$, and not doing the trade. In the case of the futures, where the price is paid at maturity, the indifference price is the solution $\alpha$ to the equation
@@
For the call, put, and swaption, where the price is received at the time of purchase, we use the SOFR overnight roll-over account as a numeraire. This allows for the computation of the selling price again analytically without numerical line search. Indeed, the selling price is the solution $\alpha$ to the equation
$EU\left(\bar{w}+\alpha-\frac{c}{\prod_{t=t_{0}}^{t_{1}-1}(1+r_{t}\delta)}\right)=EU\left(\bar{w}\right).$ (24)
With the exponential utility, we obtain
$\alpha=\frac{1}{\rho}\ln E\exp\left\{\frac{\rho c}{\prod_{t=t_{0}}^{t_{1}-1}(1+r_{t}\delta)}\right\}.$ (25)
The indifference price $\alpha$ in (23) and (25) can be computed by numerically evaluating the expectations. In the numerical computations below, we will use Monte Carlo with antithetic sampling. That is, we simulate a random sample of $N$ scenarios and obtain another $N$ scenarios by reflecting the underlying Gaussian process.
Figures 12 and 13 plot the indifference selling prices for the four SOFR derivatives studied in Section 5.2. Figure 12 displays the indifference price of the futures as a function of the initial wealth $\bar{w}$ and the risk aversion parameter $\rho$. As $\bar{w}$ increases, the price increases. With an initial financial position of cash only, the agent is less desperate for extra cash and asks for a higher price to compensate for taking on the additional risk of selling the derivative. As $\rho$ increases, the price increases. Because their initial financial position consists only of cash, the higher the value of $\rho$ the more risk-averse the agent is. They require more money to sell the derivative. This is also observed in the plots of Figure 13 which display the swaption and option price as functions of the risk aversion parameter $\rho$.
Figure 12: Selling price of SOFR three-month futures as a function of the initial wealth $\bar{w}$ and the risk aversion parameter $\rho$ (scaled by a factor of $10^{-3}$) as of 28 August 2024.
Figure 13: Selling price of SOFR swaption and options on three-month futures as a function of the risk aversion parameter $\rho$ (scaled by a factor of $10^{-3}$) as of 28 August 2024.
@@ 6 Arbitrage
The following theorem is from [23, Theorem 3], but one should note that part g of the statement there has a typo: the symbol $Q_{n}$ should be replaced by $\bar{Q}_{n}$. An alternative proof and an extended statement can be found in [33, Remark 2.120]. The ${\cal F}_{t}$}-conditional support of a random variable $X\in L^{0}(\Omega,{\cal F},P;\mathbbm{R}^{d})$ is the smallest closed-valued ${\cal F}_{t}$-measurable mapping $\mathop{\rm supp}\nolimits_{t}X:\Omega\rightrightarrows\mathbbm{R}^{d}$ such that $X\in\mathop{\rm supp}\nolimits_{t}X$ almost surely. The convex hull and the relative interior of a set $C\subseteq\mathbbm{R}^{d}$ will be denoted by $\mathop{\rm co}C$ and $\mathop{\rm ri}C$, respectively.
@@ 6.1 Arbitrage in zero-coupon bonds
By [33, Theorem 2.43.4], the continuity of $H$ implies
@@
The claim now follows from the fact that $\mathop{\rm ri}\mathop{\rm co}\mathop{\rm cl}C=\mathop{\rm ri}\mathop{\rm co}C$ for any set $C$ in a Euclidean space; see [36, Theorem 6.3]. ∎
@@
i.e. if the conditional support of the day-ahead forward curve contains every positive forward curve. This condition is far from necessary but it could be enforced e.g. by multiplying each $F_{t}(s)$ by $e^{\epsilon_{t}}$, where $\epsilon_{t}$ are independent Gaussian variables with nonzero variance. Such additional randomness could be interpreted as observation errors, as is common in the theory of affine term structure models; see e.g. [10, Section 4]. Taking the variance of $\epsilon_{t}$ sufficiently close to zero, the additional randomness would not have a material effect on portfolio optimization or indifference pricing in the face of bid-ask spreads. More importantly, one cannot trade perfectly liquid zero-coupon bonds with arbitrary maturities in practice, so the no-arbitrage conditions derived above aren’t all that relevant in real-life applications.
@@ 6.2 Arbitrage in futures market
By [33, Theorem 2.43.4], continuity of the exponential function implies
@@
The claim now follows from the fact that $\mathop{\rm ri}\mathop{\rm co}\mathop{\rm cl}C=\mathop{\rm ri}\mathop{\rm co}C$ for any set $C$ in a Euclidean space; see [36, Theorem 6.3]. ∎
@@
7 Conclusion
This paper presents a statistical SOFR term structure model, which is easily calibrated to historical data, current market conditions as well as users’ views. The model operates on a daily frequency, and it is consistent with the discrete jumps often observed on FMOC meeting dates. The model allows for fast simulations, which are needed in e.g. in risk management, portfolio optimization and indifference pricing of interest rate derivatives.
The calibration of the model breaks down into three separate steps summarized in Algorithms 1 to 3 below. In the first step, Algorithm 1 fits a parameterized forward curve to observed daily snapshots of futures quotes. Algorithms 2 and 3 then calibrate a macroeconomic model and a parametric forward curve model, respectively, to historical data. Algorithm 4 summarizes the simulation procedure once the model has been calibrated.
The overall modelling approach is modular and quite flexible. For example, the piecewise linear basis functions used to parameterize the forward curve in the numerical computations in this paper could easily be replaced by other functions. One could use e.g. the classical Nelson-Siegel three-factor parameterization. Also, the Gaussian processes used to model the macroeconomic factors and the term structure risk factors could readily be replaced by other specifications.
Instead of the macroeconomic model built in Algorithm 2, one could equally well plug in an alternative model that provides a description of the future development of the Fed rate $L$ that is used in step 6 of Algorithm 4 to transform the simulated process $x$ to the forward curve risk factors $\xi$. In particular, it would be interesting to study models where the forward curve has predictive power over the FOMC’s future interest rate decisions. Such features could be incorporated in the framework of Section 4 by regressing the macroeconomic variables $y$ on the forward curve risk factors $x$. Another interesting modification would be to regress the increments $\Delta y_{t}$ of the macroeconomic factors on their past values. Such a modification could add ”inertia” to the macroeconomic variables to describe, for example, the tendency of the Fed to move interest rates in the same direction on consecutive FOMC meeting dates.
Input: Daily time series for SOFR futures prices
1
1exOutput: Daily time series for term structure risk factors $\xi_{t}=(\xi_{t}^{k})_{k=1}^{K}$ (Figure 4)
2
31exChoose the $K$ tenors for the piecewise linear basis functions (15)
4 for each day $t$ do
5   Solve the calibration problem in (11)
Algorithm 1 Construct the historical time series of the term structure risk factors $\xi$ (Section 3)
Input: Monthly historical time series of the macroeconomic variables $(L,I,G)$
Input: Long-term medians of $(L,I,G)$ (see Table 11)
Input: Short-term median for process $L$ (The simulations of Section 5.1 used the last observed forward curve computed using Algorithm 1)
1
1exOutput: Autoregression matrix AA, the constant vectors $a_{t}$ and the covariance matrix $\Sigma$ in equation (20) (see Tables 68)
2
31exUse the non-linear transformation (19) to map the monthly values of $(L,I,G)$ to $y$
4 Estimate the autoregression matrix AA in (20) by applying linear regression to the three components of $y$
5 Calculate an estimate of the covariance matrix $\Sigma$ using the residuals
Calculate the constant vectors $a_{t}$ in equation (20) using the procedure of Section 4.2 of [27] (see Section 4.5)
Algorithm 2 Calibrate the macroeconomic model (Sections 4.2,  4.3,  4.5)
Input: Historical time series of the daily values of the term structure risk factors $\xi$ computed using Algorithm 1
Input: Historical time series of monthly values of $L$
Input: Long term medians of the forward curve risk factors $\xi$ (see Table 12)
1
1exOutput: Autoregression matrix AA, the constant vectors $a_{t}$ and the covariance matrix $\Sigma$ in equation (21) (see Tables 910)
2
31exUse the nonlinear transformation given by (17) and (18) to map the daily values of $(\xi^{k})_{k=1}^{K}$ to $x$
4 Estimate the autoregression matrix AA in (21) by applying linear regression to the $K$ components of the time series $x$
5 Calculate an estimate of the covariance matrix $\Sigma$ using the residuals
Calculate the constant vectors $a_{t}$ in equation (21) using the procedure of Section 4.2 of [27] (see Section 4.5)
Algorithm 3 Calibrate the forward curve model (Sections 4.2,  4.4,  4.5)
Input: Autoregression matrices AA, constant vectors $a_{t}$ and the covariance matrices $\Sigma$ from Algorithms 2 and 3
Input: Starting values of $(L,I,G)$ and $\xi$
Input: Start and end dates of the simulation
Input: Number of scenarios to be simulated
1
1exOutput: Simulated time series for the macroeconomic variables $(L,I,G)$ (Figure 8)
Output: Simulated time series of the forward curve $F$
2
31exSimulate Gaussian innovations $\epsilon_{t}$ of the macoeconomic model
4 Calculate scenarios of the macroeconomic risk factors $y$ using equation (20)
5 Apply the inverse of the transformations (19) to $y$ to get the scenarios for $(L,I,G)$
6 Simulate Gaussian innovations $\epsilon_{t}$ of the forward curve model
7 Calculate scenarios of the forward curve model $x$ using equation (21)
8 Apply the inverse of the transformations (17) and (18) to $x$ using simulated values of $L$ to get the scenarios for $\xi$
9 Compute scenarios of the forward curve $F$ using expression (16)
Algorithm 4 Simulate the macroeconomic variables and the SOFR term structure (Section 5.1)